OpenStudy (anonymous):

Which of the following is a polynomial with roots 5, −6, and 7? f(x) = x3 − 6x2 − 30x + 37 f(x) = x3 − 30x2 − 37x + 42 f(x) = x3 − 6x2 − 37x + 210 f(x) = x3 − 30x2 − 42x + 210 medals awarded to answer

OpenStudy (anonymous):

expand \((x-5)(x+6)(x-7)\)

OpenStudy (alekos):

That's a good start!

OpenStudy (anonymous):

grrr f(x) = x3 − 6x2 − 30x + 37?

OpenStudy (alekos):

no way. the expansion will give you one of the choices

OpenStudy (alekos):

try again

OpenStudy (anonymous):

I see the x^3 30x^2

OpenStudy (anonymous):

42

OpenStudy (anonymous):

f(x) = x3 − 30x2 − 42x + 210

OpenStudy (anonymous):

start by foiling 2 of them: \((x-5)(x+6)(x-7)=(x^2++x-30)(x-7)\) now distribute futher

OpenStudy (anonymous):

Has to be f(x) = x3 − 6x2 − 37x + 210

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thankssss

OpenStudy (anonymous):

Can you give a medal to each other? i feel guilty

OpenStudy (anonymous):

give them to the other guy, this isn't my main account

OpenStudy (anonymous):

Give me main so I can fan

OpenStudy (anonymous):

http://openstudy.com/users/bibby

OpenStudy (alekos):

finally got it!